ganeshie8 (ganeshie8):

3f(x-2) + f(2-x) = 5x f(x) = ?

OpenStudy (freckles):

\[\text{ is \it } f(x)=\frac{5}{2}(x+2)\] checking.. one sec

ganeshie8 (ganeshie8):

thats super fast xD how did you..

OpenStudy (freckles):

well the sum of odd functions should be odd

OpenStudy (freckles):

f(2-x)=f(-(x-2))=-f(x-2)

Nnesha (nnesha):

^_^ *

OpenStudy (freckles):

\[2f(x-2)=5x\]

OpenStudy (freckles):

\[f(x-2)=\frac{5x}{2} \\ \text{ \let } u=x-2 \\ \text{ then } u+2=x \\ f(u)=\frac{5}{2}(u+2)\]

OpenStudy (freckles):

checking: \[3f(x-2)+f(2-x) \\ =3\frac{5}{2}(x-2+2)+\frac{5}{2}(2-x+2) \\ =\frac{15}{2}(x)+\frac{5}{2}(4)-\frac{5}{2}x \\ =\frac{10}{2}x+10 \\ =5x+10\] but I think I made a mistake

ganeshie8 (ganeshie8):

i misplaced the solution, one sec...

OpenStudy (freckles):

I guess odd sum doesn't have to mean both addends are odd

OpenStudy (kainui):

Well I was just about to ask lol.

ganeshie8 (ganeshie8):

I see.. but that looks new and interesting... never thought about doing that before xD

ganeshie8 (ganeshie8):

so the question is still open..

OpenStudy (kainui):

I might have a solution let me finish solving this matrix...

ganeshie8 (ganeshie8):

okie :) for solving, i think we can assume f(x) is continuous/differentiable/smooth/good and a nice function...

OpenStudy (kainui):

let's do the sensible thing here and x-2=y so we have: \[3f(y)+f(-y)=5(y+2)\] Here's my trick, exploit the fact that every function is the sum of an even and odd function: \[f(x) = \frac{f(x)+f(-x)}{2}+\frac{f(x)-f(-x)}{2}=h(x)+g(x)\] h(x) is that first, even one, and g(x) is the second, odd one, pretty straight forward so far I hope? Ok now let's substitute in f(x)=h(x)+g(x) to get: \[3h(y)+3g(y)+h(-y)+g(-y)=5(y+2)\] use the properties of even and odd functions to get the first equation below and then let y be -y to get the second equation: \[4h(y)+2g(y)=5(y+2) \\ 4h(y)-2g(y)=5(-y+2)\] I don't know why I used a matrix to solve this, when it's so obvious to solve by adding and subtracting the two equations, to get: \[8h(y)=20 \\ 4g(y) = 10y\] Now plug them in since f(x)=h(x)+g(x)... Also I just realized I don't know my alphabet, it should be fgh not fhg... uh... whoops lol.

Miracrown (miracrown):

O.O

Miracrown (miracrown):

Is f(x) a linear function? if so you can begin the set up with f(x) = mx +b then we can solve for m and b

ganeshie8 (ganeshie8):

it could be linear or exponential, we won't know until we solve it.. im still going through Kainui's solution xD

OpenStudy (kainui):

I'm getting that f(x)=5/2(x-1) I think, unless I made some algebra mistake, I'm checking to make sure it meets the question's requirements.

Miracrown (miracrown):

I'm checking it too

ganeshie8 (ganeshie8):

\[8h(y)=20 \\ 4g(y) = 10y\] gives \[f(x) = \frac{5}{2}(x+1)\] right ?

OpenStudy (freckles):

and if so it is correct awesome work @Kainui

Miracrown (miracrown):

When I plugged that into the initial formula this didn't work. What did you get? For example 3f(x-2) will become 3(5/2((x-2)-1) right?

Miracrown (miracrown):

and f(2-x) is 5/2((2-x)-1)

Miracrown (miracrown):

if we add them we need 5x

OpenStudy (freckles):

\[3 \cdot \frac{5}{2}(1+x-2)+\frac{5}{2}(1+2-x) \\ \frac{15}{2}(x-1)+\frac{5}{2}(3-x) \\ \frac{15}{2}x-\frac{15}{2}+\frac{15}{2}-\frac{5}{2}x \\ \frac{10}{2}x \\ 5x\]

ganeshie8 (ganeshie8):

this f(x) = odd + even is really a neat thing !

OpenStudy (kainui):

Yeah I don't know how I figured that out, I just got lucky.

OpenStudy (kainui):

But it might be a pretty general method for solving these things.

OpenStudy (kainui):

also @freckles I don't know how you got the algebra to work out when me and mira couldn't cause I'm still struggling somewhere.

ganeshie8 (ganeshie8):

i had a solution using taylor series but im bit embarassed now to post after seeing that cute method lol

OpenStudy (freckles):

@Kainui did you use the corrected function?

Miracrown (miracrown):

Kainui sometimes I think you have a portion of Einstein's intelligence implemented somewhere in your brain

OpenStudy (freckles):

\[f(x) = \frac{5}{2}(x+1) \]

OpenStudy (kainui):

@ganeshie8 Wait show it! I was trying to find the taylor series type solution with some kind of recursion but I couldn't so that's why I did this junk! XD

OpenStudy (kainui):

Hahaha thanks @Miracrown !!! I wish I could put that on my resume right now.

OpenStudy (freckles):

don't be embarrassed I posted something totally wrong and I don't care

Miracrown (miracrown):

Kainui you can, there is copy and paste for a reason

OpenStudy (freckles):

also wasn't trying to imply what you are about to do is wrong i was saying who can be more embarrassed than me and i'm not embarrassed so you shouldn't be either :p

Miracrown (miracrown):

lol freckles let me remind kainui of this quote: ''Be who you are and say what you feel, because those who mind don't matter and those who matter don't mind''

ganeshie8 (ganeshie8):

\(3f(x-2) + f(2-x) = 5x\) plugin \(x = 2\), we get \(f(0) = 5/2\) differentiating we get : \(3f'(x-2) -f'(2-x) = 5 \implies f'(0) = 5/2\) and \(f^n(0) = 0\) for \(n\gt 1\) \(\cdots\)

OpenStudy (freckles):

nope figured it out

OpenStudy (freckles):

you used that little equation there

Miracrown (miracrown):

Here's what I did and got a similar result: 3f(x-2) + f(2-x) = 5x F(x) = mx+b 3m(x-2) + 3b+m(2-x)+b=5x 3mx-6m+3b+2m-mx+b=5x 2mx=5x and 4b-4m = 0 m = 5/2 b = 5/2 f= 5/2 (x+1)

ganeshie8 (ganeshie8):

i think that works perfectly too as it gives the function that satisfies the given equation xD freckels : \(3f'(x-2) -f'(2-x) = 5~; ~\color{Red}{\text{ plugin x = 2}~}\implies f'(0) = 5/2 \)

ganeshie8 (ganeshie8):

Oh you figured it out already okieee

OpenStudy (freckles):

I like your method @ganeshie8 and @Miracrown

OpenStudy (freckles):

and @Kainui god don't want to leave anyone out

OpenStudy (freckles):

or gosh if you prefer gosh

ganeshie8 (ganeshie8):

lol i wish to remember writing a function as "odd+even" while solving other problems xD

Miracrown (miracrown):

That's a good approach, we just need to make sure we get the right answer when using it

OpenStudy (kainui):

\(\color{blue}{\text{Originally Posted by}}\) @freckles or gosh if you prefer gosh \(\color{blue}{\text{End of Quote}}\) I prefer Jambalaya actually...

Miracrown (miracrown):

ich bin gott!!!

OpenStudy (freckles):

lol

OpenStudy (kainui):

Here's another thing I considered using that's sort of similar to the power series method @ganeshie8 the derivative of an even function is an odd, and odd is even: \[\large \left[ \frac{f(x)+f(-x)}{2} \right]' = \frac{g(x)-g(-x)}{2}\]